A $50,000 loan is taken out on a new car with the terms 12% APR (monthly) for 48 months. How much are the required monthly payments on this loan?

a.

$1,316.69

b.

$1,711.70

c.

$1,301.68

d.

$1,580.03

e.

$1,448.36
Answer #1

This question requires application of PV of annuity,

PV =

[1-(4+5)-) [1-(1+r)-n LT P= Periodic Payment r=rate per period n = number of periods

PV = $50000, n = 48 months, r = 12%/12 = 1% (monthly)

50000 =P*,1-(1+0.01)-48 50000 = P* 0.01

50000 = P * 37.9740

P = $1,316.69